OpenStudy (anonymous):

PLEASE HELP!!! If csc theta= -5/3 and theta has its terminal side in quadrant III, find the exact value of tan2 theta

OpenStudy (anonymous):

draw a triangle it is the mother of all right triangles, the 3 - 4-5 right triangle from that you can get sine, cosine everything

OpenStudy (anonymous):

like this?

OpenStudy (anonymous):

I got tan(theta) which is -3/4

OpenStudy (anonymous):

I'm not sure how tan^2 actually works...

OpenStudy (anonymous):

ok you got \(\tan(\theta)\) what are you looking for ?

OpenStudy (anonymous):

for tan^2(theta)

OpenStudy (anonymous):

are you looking for \[\tan(2\theta)\] or \[\tan^2(\theta)\]

OpenStudy (anonymous):

i am 100% sure you can square \(-\frac{3}{4}\)

OpenStudy (anonymous):

oh wait. it's the first one.

OpenStudy (anonymous):

tan2theta. my bad.

OpenStudy (anonymous):

\[\tan 2 \theta\]

OpenStudy (anonymous):

i never remember that one is it \[\frac{\tan(x)}{1-\tan^2(x)}\]

OpenStudy (anonymous):

oh no \[\frac{2\tan(x)}{1-\tan^2(x)}\]

OpenStudy (anonymous):

I'm not quite sure..

OpenStudy (anonymous):

yeah it is the second one in any case you have the number you need, it is arithmetic from here on in

OpenStudy (anonymous):

So I fill x with -3/4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got 2tan(-3/4)/1-tan^2(-3/4) = 2.7

OpenStudy (anonymous):

ooh NNOOO

OpenStudy (anonymous):

lol what?

OpenStudy (anonymous):

it is not the tangent OF \(-\frac{3}{4}\) the tangent IS \(-\frac{3}{4}\)

OpenStudy (anonymous):

yeah. but I have to find tangent 4 theta not tangent theta

OpenStudy (anonymous):

slow

OpenStudy (anonymous):

I know tangent theta = -3/4

OpenStudy (anonymous):

haha who , me, you or both?

OpenStudy (anonymous):

\[\tan(2\theta)=\frac{2\tan(\theta)}{1-\tan^2(\theta)}\]

OpenStudy (anonymous):

in this case \[\tan(\theta)=-\frac{3}{4}\]

OpenStudy (anonymous):

yeah tan2 theta. I keep saying 4 >,<

OpenStudy (anonymous):

therefore \[\tan(2\theta)=\frac{2\left(-\frac{3}{4}\right)}{1-\left(-\frac{3}{4}\right)^2}\]

OpenStudy (anonymous):

you don't take the tangent OF \(-\frac{3}{4}\) the tangent IS \(-\frac{3}{4}\)

OpenStudy (anonymous):

i see i made a mistake you said replace x by -3/4 and i said yes i should have said HELL NO you replace \(\tan(x)\) by -3/4

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

do I simplify the last equation?

OpenStudy (anonymous):

lol it is not an equation it is a number and yes, no one wants to look at \[\frac{2\left(-\frac{3}{4}\right)}{1-\left(-\frac{3}{4}\right)^2}\] it is just some fraction

OpenStudy (anonymous):

I got -24/7

OpenStudy (anonymous):

good it is right there is another famous right triangle with one leg 7 and the other 24

OpenStudy (anonymous):

You are very helpiful. Thanks so much :) If you don't mind, would you help me with just one more?

OpenStudy (anonymous):

sure hope i don't mess up this time

OpenStudy (anonymous):

if 90 degrees < theta < 180 degrees and cos theta=-4/5, find sin 4 theta

OpenStudy (anonymous):

crap

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

is it complicated?

OpenStudy (anonymous):

\[\sin(\theta)=\frac{3}{5}\]

OpenStudy (anonymous):

\[\sin(2\theta)=2\cos(\theta)\sin(\theta)\] plug in the numbers let me know what you get

OpenStudy (anonymous):

Do I plug in the degrees?

OpenStudy (anonymous):

oh no it is like last time plug in the numbers you know

OpenStudy (anonymous):

\[2\times \frac{-3}{5}\times \frac{4}{5}\]

OpenStudy (anonymous):

-24/25?

OpenStudy (anonymous):

ok so that is \[\sin(2\theta)\]

OpenStudy (anonymous):

ooh now I need to get sin4theta

OpenStudy (anonymous):

you want \[\sin(4\theta)\] so we repeat the process, but this time with \[\sin(2\theta)=\frac{-24}{25}\]

OpenStudy (anonymous):

sin4theta = 4 * -24/25 * -4/5 ?

OpenStudy (anonymous):

\[\sin(4\theta)=2\cos(2\theta)\sin(2\theta)\]

OpenStudy (anonymous):

or do I eliminate the negatives?

OpenStudy (anonymous):

we have \[\sin(2\theta)\] we need \[\cos(2\theta)\]

OpenStudy (anonymous):

ooOh

OpenStudy (anonymous):

it is not -4/5 we need the cosine of two theta

OpenStudy (anonymous):

would it be? 2sintheta * cos theta

OpenStudy (anonymous):

\[\cos(2x)=2\cos^2(x)-1\] so we have to find that as well

OpenStudy (anonymous):

do I plug in -24/25?

OpenStudy (anonymous):

\[\cos(2x)=2(\frac{16}{25})-1\]

OpenStudy (anonymous):

\[\cos(2x)=\frac{7}{25}\]

OpenStudy (anonymous):

now plug that in

OpenStudy (anonymous):

4 * -24/25 *7/25 ?

OpenStudy (anonymous):

no not 4, 2

OpenStudy (anonymous):

same formula, different sine and cosine

OpenStudy (anonymous):

just change your 4 to a 2

OpenStudy (anonymous):

i got -336/625

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

i calculated 2 * -24/25 *7/25

OpenStudy (anonymous):

then it is right

OpenStudy (anonymous):

THANK YOU SO MUCH. Sorry for taking so much of your time!

OpenStudy (anonymous):

no problem yw